# Krylov Distribution

### v1

@@ I Introduction
Most existing applications of Krylov-space methods focus on unitary time evolution. In this setting, the spreading of a time-evolved state $|\psi(t)\rangle=e^{-iHt}|\psi_{0}\rangle$ along the Krylov chain is quantified by the Krylov complexity Parker et al. (2019); Rabinovici et al. (2021); Balasubramanian et al. (2022); Caputa et al. (2022); Barbón et al. (2019); Alishahiha and Banerjee (2023). Krylov complexity has proven to be a sensitive diagnostic of operator growth, quantum chaos, and universality across a wide range of quantum systems, including spin chains, random matrix models, conformal field theories, and holographic setups, as reviewed in Refs. Nandy et al. (2025); Rabinovici et al. (2025). These developments have clarified how dynamical complexity is governed by the growth of Lanczos coefficients and the emergent geometry of Krylov space.
@@ II Krylov Resolvent Amplitudes and the Krylov Distribution
where $\xi$ is a real spectral parameter. Acting on the reference state produces the resolvent-dressed state
@@ III Krylov Resolvent Amplitudes and the Thermodynamic Limit
The large-$n$ behavior of $Q_{n}(E)$ in the interior of the spectrum is oscillatory. Away from spectral edges, the leading Plancherel–Rotach asymptotics read
@@ VI Conclusions and Outlook
In this work, we introduced the Krylov distribution $\mathcal{D}(\xi)$ as a static, resolvent-based diagnostic that organizes inverse-energy response in terms of Krylov-space structure. By treating the resolvent-dressed state $(H-\xi)^{-1}|\psi_{0}\rangle$ as a primary object, this framework provides a physically transparent way to relate spectral properties, static response functions, and the emergent geometry of Krylov space. Combining asymptotic analysis, exact solutions in analytically tractable models, and numerical studies of interacting systems, we demonstrated that $\mathcal{D}(\xi)$ exhibits robust and interpretable scaling behavior across gapped, gapless, and critical regimes.
@@
A central result of our analysis is that the Krylov distribution sharply distinguishes whether the spectral parameter $\xi$ lies inside or outside the many-body spectrum. When $\xi$ is separated from the spectrum by a finite gap, the resolvent-dressed state is exponentially localized in Krylov space and $\mathcal{D}(\xi)$ saturates to an $\mathcal{O}(1)$ value in the thermodynamic limit. By contrast, when $\xi$ lies within a continuous part of the spectrum associated with an absolutely continuous spectral measure, resolvent amplitudes remain extended along the Krylov chain in an averaged sense over a broad range of Krylov indices, leading to extensive growth $\mathcal{D}(\xi)\sim N/2$ for an effective Krylov dimension $N$. This spectral dichotomy provides a direct and intuitive diagnostic that complements conventional probes such as spectral functions or densities of states.
@@
The exactly solvable models analyzed in this work illustrate these principles from complementary perspectives. Systems with constant Lanczos coefficients, encompassing tight-binding models and certain free-fermion systems, display a smooth interpolation of the Krylov distribution from exponential saturation outside the spectral band to linear growth inside it, capturing the essential features of bounded continuous spectra. In contrast, the displaced harmonic oscillator, characterized by an unbounded discrete spectrum and square-root-growing off-diagonal Lanczos coefficients, yields closed-form resolvent amplitudes $\psi_{n}(\xi)$ and a Krylov distribution exhibiting sharp resonant peaks at individual eigenvalues, separated by regions of strong Krylov-space localization. Finally, the $\mathrm{SU}(1,1)$ chain with linearly growing off-diagonal coefficients provides an analytically tractable model of continuous spectra with asymptotically linear Lanczos growth. In this setting, $\mathcal{D}_{N}(\xi)\sim N/\ln N$ signals power-law delocalization in Krylov space and the absence of spectral gaps. Together, these examples span a broad taxonomy of spectral structures—bounded and unbounded, discrete and continuous—and their distinct manifestations in Krylov space.
@@
Our numerical study of the mixed-field Ising chain further demonstrates that while $\mathcal{D}(\xi)$ is primarily sensitive to spectral boundaries, it can also retain signatures of integrability versus chaos for structured initial states. Product states yield larger oscillations and enhanced magnitude of $\mathcal{D}(\xi)$ in the integrable regime, whereas chaotic dynamics produce smoother and more suppressed profiles. These distinctions are partially preserved for Gibbs-like superpositions but are significantly reduced under ensemble averaging with random states. As a result, $\mathcal{D}(\xi)$ should be viewed not as a universal chaos indicator, but as a tunable probe whose diagnostic power depends on the choice of initial state and the degree of spectral averaging.
@@
Beyond its role as a spectral diagnostic, the Krylov distribution is closely connected to static response and quantum geometry. We showed that the fidelity susceptibility corresponds directly to the total weight $\sum_{n}|\psi_{n}(\xi)|^{2}$ of the resolvent-dressed state evaluated at the ground-state energy. More generally, the quantum geometric tensor admits a natural decomposition in terms of Krylov-resolved resolvent amplitudes, highlighting the Krylov basis as an efficient and physically meaningful representation for encoding quantum geometry and adiabatic response. Higher inverse-gap moments can likewise be generated through successive derivatives of $\psi_{n}(\xi)$ with respect to $\xi$, providing a systematic framework for accessing static response functions of increasing order.
@@
At quantum critical points, the Krylov distribution is expected to exhibit characteristic scaling behavior. In gapless systems where the finite-size gap scales as $\Delta\sim 1/L$, the exponential energy resolution inherent to Krylov space implies that resolving energies down to $\Delta$ requires a Krylov depth that grows logarithmically with system size. This naturally leads to $\mathcal{D}(\xi)\sim\log L$ at criticality, reflecting the slow divergence of correlation lengths typical of many quantum phase transitions. A more detailed exploration of this connection may provide a practical route for diagnosing critical behavior in numerical simulations and, potentially, in experimental platforms.
@@
Our results also clarify the relationship between static and dynamical Krylov diagnostics. While Krylov complexity $\mathcal{C}(t)$ characterizes the spreading of the time-evolved state $e^{-iHt}|\psi_{0}\rangle$ in Krylov space and is sensitive to operator growth and scrambling, the Krylov distribution $\mathcal{D}(\xi)$ encodes how resolvent-dressed states explore Krylov space as a function of energy. These two perspectives are related through the Laplace transform,
@@
which maps time-domain spreading to energy-domain resolution. Importantly, this relation does not imply that $\mathcal{D}(\xi)$ can be reconstructed from finite-time dynamics alone. The resolvent probes inverse-energy structure and spectral organization in a manner that is qualitatively distinct from dynamical diagnostics, particularly near spectral gaps, edges, and critical points.
Several directions for future work emerge naturally from this study. Quantitative measures derived from fluctuations or correlations of $\mathcal{D}(\xi)$ along the spectral axis may provide additional probes of integrability and chaos beyond traditional level statistics, and a systematic finite-size scaling analysis would be valuable in assessing their universality. Extending the present framework to many-body localized systems is another promising direction, where one expects Krylov-space localization to persist even within the spectrum. More broadly, generalizations to non-Hermitian Hamiltonians or Lindbladian dynamics could shed light on spectral and dynamical properties of open quantum systems. On the mathematical side, a more rigorous understanding of universal scaling laws—such as the $N^{2/3}$ behavior near generic spectral edges—would strengthen connections to random matrix theory and the asymptotic theory of orthogonal polynomials.
Finally, while direct experimental access to the full Krylov distribution remains challenging, its connection to fidelity susceptibility and quantum geometric tensors suggests that certain aspects of $\mathcal{D}(\xi)$ may be accessible in quantum simulators using existing measurement protocols.
In summary, we have established the Krylov distribution $\mathcal{D}(\xi)$ as a static counterpart to dynamical Krylov complexity. By linking spectral structure, scaling behavior, quantum geometry, and many-body dynamics within a unified Krylov-space framework, this work lays a foundation for systematically probing energy-resolved structure in quantum many-body systems.
@@ Appendix C Resolvent Krylov Basis and Generalized Krylov Constructions
More generally, one may consider extended Krylov constructions generated by the joint action of $H$ and $H^{-1}$ on $|\psi_{0}\rangle$, thereby interpolating between dynamical and resolvent-based descriptions. A systematic analysis of such hybrid constructions is left for future work.

### v2

@@ I Introduction
Most existing applications of Krylov-space methods focus on unitary time evolution. In this setting, the spreading of a time-evolved state $|\psi(t)\rangle=e^{-iHt}|\psi_{0}\rangle$ along the Krylov chain is quantified by the Krylov complexity Parker et al. (2019); Rabinovici et al. (2021); Balasubramanian et al. (2022); Caputa et al. (2022); Barbón et al. (2019); Alishahiha and Banerjee (2023). Krylov complexity has proven to be a sensitive diagnostic of operator growth, quantum chaos, and universality across a wide range of quantum systems, including spin chains, random matrix models, conformal field theories, and holographic setups Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Vasli et al. (2024); Imani et al. (2025); Rabinovici et al. (2022); Scialchi et al. ; Trigueros and Lin (2022); Español and Wisniacki (2023); Erdmenger et al. (2023); Huh et al. (2024); Camargo et al. (2024); Nandy et al. (2025b); Bhattacharjee and Nandy (2025); Balasubramanian et al. (2025); Baggioli et al. (2025b); Alishahiha et al. (2025); Astaneh and Vardian (2025); Bhattacharya and Jana (2024); Huh et al. (2025); Baggioli et al. (2025a), as reviewed in Refs. Nandy et al. (2025a); Rabinovici et al. (2025). These developments have clarified how dynamical complexity is governed by the growth of Lanczos coefficients and the emergent geometry of Krylov space.
@@ II Krylov Resolvent Amplitudes and the Krylov Distribution
where $\xi$ is a real spectral parameter. Acting on the reference state produces the resolvent-dressed state [^1] This framework can also be extended to resolvent-dressed operators; see Lunt et al. (2025) for a related study.
@@ III Krylov Resolvent Amplitudes and the Thermodynamic Limit
The large-$n$ behavior of $Q_{n}(E)$ in the interior of the spectrum is oscillatory. Away from spectral edges, the leading Plancherel–Rotach asymptotics read Szegö (1939); Deift (1999); Simon (2005)
@@ VI Conclusions and Outlook
In this work, we introduced the Krylov distribution $\mathcal{D}(\xi)$ as a static, resolvent-based diagnostic that organizes inverse-energy response in terms of Krylov-space structure. By elevating the resolvent-dressed state $(H-\xi)^{-1}|\psi_{0}\rangle$ to a primary object, this framework provides a transparent connection between spectral properties, static response functions, and the emergent geometry of Krylov space. Combining asymptotic analysis, exact solutions in analytically tractable models, and numerical studies of interacting systems, we showed that $\mathcal{D}(\xi)$ exhibits robust and interpretable scaling behavior across gapped, gapless, and critical regimes.
@@
A central result is that the Krylov distribution sharply distinguishes whether the spectral parameter $\xi$ lies inside or outside the many-body spectrum. When $\xi$ is separated from the spectrum by a finite gap, the resolvent-dressed state is exponentially localized in Krylov space and $\mathcal{D}(\xi)$ saturates to an $\mathcal{O}(1)$ value in the thermodynamic limit. By contrast, when $\xi$ lies within a continuous part of the spectrum associated with an absolutely continuous spectral measure, resolvent amplitudes remain extended along the Krylov chain in an averaged sense, leading to extensive growth $\mathcal{D}(\xi)\sim N/2$ for an effective Krylov dimension $N$. This spectral dichotomy provides a direct diagnostic that complements conventional probes such as spectral functions and densities of states.
@@
The exactly solvable models illustrate these principles from complementary perspectives. Systems with constant Lanczos coefficients display a smooth interpolation from exponential saturation outside the spectral band to linear growth inside it, capturing the essential features of bounded continuous spectra. The displaced harmonic oscillator, characterized by an unbounded discrete spectrum and square-root-growing coefficients, yields closed-form amplitudes $\psi_{n}(\xi)$ and a Krylov distribution with sharp resonant peaks at eigenvalues separated by strongly localized regions. The $\mathrm{SU}(1,1)$ chain with linearly growing coefficients provides a tractable model of continuous spectra with asymptotically linear Lanczos growth; in this case, $\mathcal{D}_{N}(\xi)\sim N/\ln N$ signals power-law delocalization and the absence of spectral gaps. Together, these examples span bounded and unbounded, discrete and continuous spectra and their distinct manifestations in Krylov space.
@@
Our numerical study of the mixed-field Ising chain further shows that while $\mathcal{D}(\xi)$ is primarily sensitive to spectral boundaries, it can also retain signatures of integrability versus chaos for structured initial states. Product states exhibit larger oscillations and enhanced magnitude of $\mathcal{D}(\xi)$ in the integrable regime, whereas chaotic dynamics produce smoother and more suppressed profiles. These distinctions are partially preserved for Gibbs-like superpositions but are substantially reduced under ensemble averaging with random states. Thus, $\mathcal{D}(\xi)$ should not be viewed as a universal chaos indicator, but rather as a tunable probe whose diagnostic power depends on the reference state and the degree of spectral averaging.
@@
Beyond its role as a spectral diagnostic, the Krylov distribution is closely tied to static response and quantum geometry. The fidelity susceptibility corresponds directly to the total weight $\sum_{n}|\psi_{n}(\xi)|^{2}$ of the resolvent-dressed state evaluated at the ground-state energy. More generally, the quantum geometric tensor admits a natural decomposition in terms of Krylov-resolved amplitudes, highlighting the Krylov basis as an efficient representation of quantum geometry and adiabatic response. Higher inverse-gap moments can likewise be generated through derivatives of $\psi_{n}(\xi)$ with respect to $\xi$, providing a systematic framework for static response functions of increasing order.
@@
At quantum critical points, the Krylov distribution is expected to exhibit characteristic scaling. In gapless systems with finite-size gap $\Delta\sim 1/L$, the exponential energy resolution inherent to Krylov space implies that resolving energies down to $\Delta$ requires a Krylov depth growing logarithmically with system size, naturally leading to $\mathcal{D}(\xi)\sim\log L$ at criticality. A more detailed investigation of this connection may provide a practical route to diagnosing critical behavior in numerical and experimental settings.
@@
Our results also clarify the relationship between static and dynamical Krylov diagnostics. While Krylov complexity $\mathcal{C}(t)$ tracks the spreading of $e^{-iHt}|\psi_{0}\rangle$ and is sensitive to operator growth and scrambling, the Krylov distribution $\mathcal{D}(\xi)$ encodes how resolvent-dressed states explore Krylov space as a function of energy. These perspectives are related through the Laplace transform,
@@
which connects time-domain spreading to energy-domain resolution. Nevertheless, $\mathcal{D}(\xi)$ probes inverse-energy structure in a manner qualitatively distinct from finite-time dynamical diagnostics, particularly near spectral gaps, edges, and critical points.
More broadly, given a reference state $|\psi_{0}\rangle$, the Krylov basis generated by repeated applications of $H$ provides a unified framework for analyzing general operator functions $f(H)$. Time evolution and the resolvent correspond to two specific choices, $f_{t}(H)=e^{-iHt}$ and $f_{\xi}(H)=(H-\xi)^{-1}$, but many others arise in applications, including transfer functions, exponential and trigonometric propagators, fractional powers, and the sign function (see, for example, Güttel (2017)). In general, any dressed state of the form $f(H)|\psi_{0}\rangle$ admits a Krylov expansion governed by the same Lanczos coefficients (see, e.g., François and Ravera (2025c, a, b)). The choice of $f$ selects which spectral features are emphasized, while the Krylov geometry organizes how that information is distributed.
Several directions for future work naturally emerge. Quantitative measures derived from fluctuations or correlations of $\mathcal{D}(\xi)$ along the spectral axis may provide additional probes of integrability and chaos beyond traditional level statistics, and systematic finite-size scaling would help assess their universality. Extending the framework to many-body localized systems, where Krylov-space localization may persist even within the spectrum, is another promising direction. Generalizations to non-Hermitian Hamiltonians or Lindbladian dynamics could shed light on open quantum systems. On the mathematical side, a more rigorous understanding of universal scaling laws—such as $N^{2/3}$ behavior near generic spectral edges—would strengthen connections to random matrix theory and orthogonal polynomial asymptotics.
Finally, although direct experimental access to the full Krylov distribution remains challenging, its connection to fidelity susceptibility and quantum geometric tensors suggests that certain aspects of $\mathcal{D}(\xi)$ may be accessible in quantum simulators using existing measurement protocols.
In summary, we have established the Krylov distribution $\mathcal{D}(\xi)$ as a static counterpart to dynamical Krylov complexity. By linking spectral structure, scaling behavior, quantum geometry, and many-body dynamics within a unified Krylov-space framework, this work lays the foundation for systematically probing energy-resolved structure in quantum many-body systems.
@@ Appendix C Resolvent Krylov Basis and Generalized Krylov Constructions
More generally, one may consider extended Krylov constructions generated by the joint action of $H$ and $H^{-1}$ on $|\psi_{0}\rangle$ Druskin and Knizhnerman (1998); Jagels and Reichel (2009); Daas and Gould (2025), thereby interpolating between dynamical and resolvent-based descriptions. A systematic analysis of such hybrid constructions is left for future work.